OpenStudy (anonymous):

One number is 9 times a first number. A third number is 100 more than the first number. If the sum of the three numbers is 1035, find the numbers.

OpenStudy (welshfella):

let first number = x then second number = 9x third number = x + 100 so add these 3 expressions and equate them to 1035

OpenStudy (welshfella):

then solve this equation for x

OpenStudy (anonymous):

so add them together equal to 1035?

OpenStudy (welshfella):

add the 3 expressions x + 9x + x + 100 and this equals 1035 so we have x + 9x + x + 100 = 1035

OpenStudy (anonymous):

X = -100/11

OpenStudy (welshfella):

?? what does x + 9x + x + 100 come to? remember you can only add LIKE TERMS

OpenStudy (anonymous):

11x+100?

OpenStudy (welshfella):

yes and this equals 1035 so we have 11x + 100 = 1035

OpenStudy (anonymous):

so X=85?

OpenStudy (welshfella):

right so you just work out the other 2 numbers by plugging x = 85 into 9x and x=85 into x + 100

OpenStudy (anonymous):

Thanks! I think i understand now! so that means that the numbers are 85, 738, and 185?

OpenStudy (welshfella):

well 185 is right but try 9*85 again

OpenStudy (anonymous):

765** Would i put them in any order?

OpenStudy (welshfella):

no any order and you can check to see if they are right by adding them up . They should total 1035.

OpenStudy (anonymous):

they do. Thanks for the help!

OpenStudy (welshfella):

in fact they do.

OpenStudy (welshfella):

yw